# Scalable Message Passing Neural Networks: No Need for Attention in Large Graph Representation Learning

### v1

@@ 5 Experimental Validation
Model Removed ogbn-proteins pokec ogbn-arxiv ogbn-products SGFormer [12] N/A 79.53 ± 0.38 73.76 ± 0.24 72.63 ± 0.13 89.09 ± 0.10 SMPNN N/A 83.15 ± 0.24 79.76 ± 0.19 73.75 ± 0.24 90.61 ± 0.05 SMPNN Residual 68.49 ± 2.59 68.17 ± 8.22 39.67 ± 14.60 89.69 ± 0.05 SMPNN $\alpha$ 82.90 ± 0.34 80.10 ± 0.12 73.00 ± 0.59 90.77 ± 0.04 SMPNN FF 80.51 ± 0.61 78.40 ± 0.20 73.25 ± 0.58 90.01 ± 0.10 SMPNN GCN LayerNorm 80.74 ± 0.91 79.42 ± 0.15 74.46 ± 0.22 90.54 ± 0.08
@@
Deep Models are Possible with SMPNNs. It is well known that conventional message-passing GNNs [30, 23, 47] are restricted to shallow architectures, since their performance degrades when stacking many layers. Next, we demonstrate that deep models are possible with SMPNNs. In particular, we perform an experiment in which we progressively increase the number of SMPNN layers while keeping all other hyperparameters constant for the ogbn-arxiv and ogbn-proteins datasets, from 2 to 12 layers. Adding up to 6 layers seems to improve performance, and it plateaus thereafter. Additionally, we perform another experiment in which we follow the same procedure but for an SMPNN without residual connections after convolutions. In this case, we observe a clear drop in performance after 4 layers, which aligns with our theoretical understanding from Section 4.
@@
GPU Memory Scaling. We proceed to examine the model’s scalability utilizing the ogbn-products dataset and randomly sampling a subset of nodes for graph mini-batching (from 10K to 100K), following the analysis in [12]. In Figure 2, we report the maximum GPU memory usage in GB against the number of edges in the subgraph induced by the sampled nodes. We can see that it asymptotes towards linearity in the number of edges as discussed in Section 3.3. For small subgraphs $N$ dominates the complexity term and results in a steeper slope. Additionally, we compare the computational overhead of different SMPNN configurations to that of SGFormer on a Tesla V100 GPU, see Figure 3. Interestingly, the scaling factors $\alpha$ seem to have a substantial computational impact as the number of nodes increases, which, combined with the ablations in Table 6, suggests it may be better to remove it for very large graphs. Moreover, we find that the GPU consumption of SMPNNs can be substantially reduced when omitting the pointwise feedforward part of the block. This reduces computation cost substantially below that of SGFormer. Hence, one can trade GPU usage for performance: note that as shown in Table 6, SMPNNs without feedforwards still outperform the SGFormer baseline. SGFormer does not use feedforward networks or traditional attention mechanisms. Instead, it employs a simplified linear version of attention and even omits the softmax activation function.
@@ 6 Conclusion
Our framework, Scalable Message Passing Neural Networks (SMPNNs), overcomes scalability issues in traditional GNNs and is based on best practices found in the LLM literature. In particular, it allows for deep architectures without degradation in performance, which is a common limitation of GNNs. SMPNNs scale efficiently to large graphs, outperforming Graph Transformers without using global attention. The lack of a need for attention can be attributed to the local nature of most large graph transductive learning tasks. Empirical validation across diverse datasets confirms SMPNNs’ competitive performance and supports the theoretical justification of residual connections. This work opens new possibilities for leveraging deep architectures in large-scale graph learning.
Acknowledgements
HSOB would like to thank Qitian Wu for the early discussions on Graph Transformers. AK acknowledges financial support from an NSERC Discovery Grant No. RGPIN-2023-04482 and No. DGECR-2023-00230. AK also acknowledges that resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute.
Appendix A Augmenting SMPNNs with Attention
@@
Linear Global Attention. Global attention can be seen as computing message passing over a fully-connected dense graph. In our block, attention is computed with respect to a virtual node, instead of the full $\mathcal{O}(N^{2})$ attention. Linear attention is more suitable for scaling to large graphs. Given the input query, key, and value matrices, $\mathbf{Q}=\mathbf{X}^{(l)}\mathbf{W}_{Q}\in\mathbb{R}^{N\times D}$, $\mathbf{K}=\mathbf{X}^{(l)}\mathbf{W}_{K}\in\mathbb{R}^{N\times D}$, $\mathbf{V}=\mathbf{X}^{(l)}\mathbf{W}_{V}\in\mathbb{R}^{N\times D}$ we apply the following operations. First, we compute a summed query value considering the contribution from all nodes, and normalize both the query and key matrices:
$\mathbf{Q}_{\text{sn}}=\frac{\sum_{i=1}^{N}\mathbf{Q}_{i}}{||\sum_{i=1}^{N}\mathbf{Q}_{i}||_{2}}\in\mathbb{R}^{D},\,\,\mathbf{K}_{n}=\frac{\mathbf{K}}{\|\mathbf{K}\|_{2}}\in\mathbb{R}^{N\times D}.$ (13)
Based on these we can compute attention
$\mathbf{A}_{attention}=\text{softmax}(\mathbf{Q}_{\text{sn}}\mathbf{K}_{n}^{T})\in\mathbb{R}^{1\times N},$ (14)
and use it to aggregate the values
$\mathbf{X}^{(l+1)}=\left(\mathbf{A}_{attention}\otimes\mathbf{1}_{N}\right)\mathbf{V}\in\mathbb{R}^{N\times D},$ (15)
where $\mathbf{1}_{N}$ is a row vector of ones with dimension $1\times N$ and $\otimes$ is the Kronecker product. The operation $\mathbf{A}_{attention}\otimes\mathbf{1}_{N}$ results in and $N\times N$ matrix, with all rows sharing the same attention coefficients, which is later multiplied by the value matrix $\mathbf{V}$ to obtain the final updated node features $\mathbf{X}^{(l+1)}$. The operations above can be extended to encompass multi-head attention. In such case we would have a total of $H$ attention heads and to compute the final global feature matrix we aggregate the contribution from all heads:
$\mathbf{X}^{(l+1)}=\sum_{h=1}^{H}\left(\mathbf{A}_{attention}^{(h)}\otimes\mathbf{1}_{N}\right)\mathbf{V}^{(h)}\in\mathbb{R}^{N\times D}.\ {-0.1cm}$ (16)
@@
The update equations above are reminiscent of hybrid Graph Transformers [14, 10, 11, 12], and we find that they do not provide tangible improvements over standard SMPNNs. Note that we also experimented with other attention mechanisms, which did not help.
Appendix B Proofs
This appendix contains all the proofs for our theoretical guarantees discussed in Section 4.2.
B.1 Proofs for Universality
The proofs of both Theorems 4.1 and 4.2 revolve around verifying the conditions (or their failure) stated in [48, Theorem 3.4]. This result characterizes situations where a continuous transformation of the inputs to a universal approximator preserves its universal approximation property. Importantly, when the output space is Euclidean (e.g., $\mathbb{R}$ in our case), it is both necessary and sufficient for the transformation applied to the input of the model class to be injective. Consequently, our proofs for both results primarily focus on the injectivity (or lack thereof) of the feature transformations in equation 11 and equation 12, respectively
In what follows, for any activation function $\zeta:\mathbb{R}\to\mathbb{R}$ we use $\mathcal{MLP}_{N\times D}^{\zeta}$ to denote the class of MLPs with activation function $\zeta$, mapping $\mathbb{R}^{N\times D}$ to $\mathbb{R}$. Note that this is mildly more general than the class $\mathcal{MLP}_{N\times D}$ considered in our main text, which used the specification $\zeta=\operatorname{SiLU}$. We consider the following regularity condition on $\zeta$.
Assumption B.1 (Regular Activation).
The activation function $\zeta:\mathbb{R}\to\mathbb{R}$ is continuous and either:
  1. (i)
    [49]: $\zeta$ is non-polynomial,
  2. (ii)
    [42]: $\zeta$ is non-affine and there exists some $t\in\mathbb{R}$ at which $\zeta$ is continuously differentiable and $\zeta^{\prime}(t)\neq 0$.
B.1.1 No universality without the residual connection
We first show that adding a single normalized graph convolution layer preceding an MLP is enough to negate the universal approximation property of the composite model in equation 11.
Lemma B.2 (Loss of Injectivity).
Let $N,D$ be positive integers with $N\geq 2$. Let $\mathcal{G}$ be a complete graph on $N$ nodes, let $\tilde{\mathbf{A}}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(I_{i\sim j\vee i=j}/\sqrt{\operatorname{deg}(v_{i})\operatorname{deg}(v_{j})})_{i,j=1}^{N}$, and let $\mathbf{W}\in\mathbb{R}^{D\times D}$. Then, the map
$\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}:\mathbb{R}^{N\times D}$ $\to\mathbb{R}^{N\times D}$
$\mathbf{X}$ $\mapsto\tilde{\mathbf{A}}\mathbf{X}\mathbf{W}$
is not injective.
Proof of Lemma B.2.
@@ B.1.1 No universality without the residual connection
The following is a more technical version of Theorem 4.1.
Lemma B.3 (No Universality without Residual Connection).
Let $\zeta$ satisfy Assumption B.1. Let $N,D$ be positive integers with $N\geq 2$. Let $\mathcal{G}$ be a complete graph on $N$ nodes, $\mathbf{W}\in\mathbb{R}^{D\times D}$, and let $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}$ be as in Lemma B.2. The class of maps $f:\mathbb{R}^{N\times D}\to\mathbb{R}$ with representation
$f=\operatorname{MLP}\circ\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}$
where $\operatorname{MLP}:\mathbb{R}^{N\times D}\to\mathbb{R}$ is an MLP with activation function $\zeta$, is not dense for the topology of uniform convergence on compacts sets in $\mathcal{C}(\mathbb{R}^{N\times D},\mathbb{R})$.
Proof.
By [49, Theorem 3.1] and [42, Theorem 3.2] the class $\mathcal{MLP}_{N\times D}^{\zeta}$ is dense in $\mathcal{C}(\mathbb{R}^{N\times D},\mathbb{R})$ for the topology of uniform convergence on compacts sets since $\zeta$ was assumed to be nonpolynomial [^1] Note that this can be relaxed further since we are using deep, and not shallow, MLPs by [42, Theorem 3.2]. However, this is not the main point of our analysis.. Since $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}$ is continuous then [48, Proposition 3.7] implies that the class
$\mathcal{MLP}_{N\times D}^{\zeta}\circ\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big{\{}\operatorname{MLP}\circ\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}:\,\operatorname{MLP}\in\mathcal{MLP}_{N\times D}^{\zeta}\big{\}},$
is dense in $\mathcal{C}(\mathbb{R}^{N\times D},\mathbb{R})$ for the topology of uniform convergence on compacts sets if and only if $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}$ is injective. By Lemma equation B.2, this map is never injective whenever $N>1$. Thus, $\mathcal{MLP}_{N\times D}^{\zeta}\circ\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}$ is not dense in $\mathcal{C}(\mathbb{R}^{N\times D},\mathbb{R})$ for the topology of uniform convergence on compacts sets; i.e. $\mathcal{MLP}_{N\times D}^{\zeta}\circ\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}$ is not a universal approximator in the sense of Definition 4.3. ∎
Next, we show that adding the residual connection recovers the universal approximation property.
B.1.2 Universality with the residual connection
We will consider the case reflective of the networks at initialization, where the graph convolution layer with residual connection is initialized randomly. We consider weights with i.i.d. random Gaussian initializations with appropriately scaled variance.
Lemma B.4 (Injectivity Regained).
Let $N,D$ be positive integers. Fix a hyperparameter $\varsigma>0$. Let $\mathcal{G}$ be a complete graph on $N$ nodes, let $\tilde{\mathbf{A}}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(I_{i\sim j\vee i=j}/\sqrt{\operatorname{deg}(v_{i})\operatorname{deg}(v_{j})})_{i,j=1}^{N}$, and let $\mathbf{W}$ be a $D\times D$ random matrix with i.i.d. entries $W_{i,j}\sim N(0,\varsigma^{2})$. If $\varsigma<1/(9D^{3/2})$ then
$\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}:\mathbb{R}^{N\times D}$ $\to\mathbb{R}^{N\times D}$
$\mathbf{X}$ $\mapsto\tilde{\mathbf{A}}\mathbf{X}\mathbf{W}+\mathbf{X}$
is injective with probability at-least $1-e^{-D/2}$.
Our proof of the positive counterpart to Lemma B.3, namely Lemma B.4, will rely on the following result from random matrix theory.
Lemma B.5 (Concentration Version of Gordon’s Theorem: Corollary 5.35 in [50]).
Let $\mathbf{A}$ be an $N\times n$ matrix whose entries are independent standard normal random variables. For every $t\geq 0$, the following holds
$\sqrt{N}-\sqrt{n}-t\leq s_{\min}(\mathbf{A})\leq s_{\max}(\mathbf{A})\leq\sqrt{N}+\sqrt{n}+t$ (22)
with probability at least $1-e^{-t^{2}/2}$; where $s_{\min}(\mathbf{A})$ and $s_{\max}(\mathbf{A})$ are respectively the smallest and largest singular values of the matrix $\mathbf{A}$.
Proof.
Step 1 - Reduction of injectivity problem to eigenvalue computation
Since $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}$ is a linear map then the dimension theorem/rank-nullity theorem/splitting lemma implies that is injective if and only if its kernel only contains the zero vector; i.e.
$\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}(\mathbf{X})=0\Leftrightarrow\mathbf{X}=0.$ (23)
Using the Kronecker product $\otimes$, see [51, Section 4.2], and the “flattenning” vectorization map $\operatorname{vec}:\mathbb{R}^{N\times D}\mapsto\mathbb{R}^{ND}$ which sends any $N\times D$ matrix to the vector obtained from ordering its entries in lexicographic order (according to their indices), see [51, Definition 4.2.9], we may rewrite
$(I_{ND}+{\tilde{\mathbf{A}}}\otimes\mathbf{W})\,\operatorname{vec}(\mathbf{X})=\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}(\mathbf{X})=0\Leftrightarrow\mathbf{X}=0,$ (24)
where $I_{ND}$ is the $ND\times ND$-dimensional identity matrix. A necessary condition in equation 24 is that $I_{ND}+{\tilde{\mathbf{A}}}\otimes\mathbf{W}$ has to be invertible. Let $T\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}-{\tilde{\mathbf{A}}}\otimes\mathbf{W}$; then, $I_{ND}+{\tilde{\mathbf{A}}}\otimes\mathbf{W}=I_{ND}-T$. The linear operator $I_{ND}-T$ is invertible  [^2] See [52, Theorem 1] for a complete characterization of the convergence of Neumann series, even in the infinite dimensional setting. if $\|T\|_{2}<1$. Since $T=-{\tilde{\mathbf{A}}}\otimes\mathbf{W}$, $\|-T\|_{2}=\|T\|_{2}$ then $(I_{ND}+{\tilde{\mathbf{A}}}\otimes\mathbf{W})$ is injective if
$\|{\tilde{\mathbf{A}}}\otimes\mathbf{W}\|_{2}<1.$ (25)
Recall that, $\|\cdot\|_{2}$ is the square-root of the sum of the squared eigenvalues of the matrix ${\tilde{\mathbf{A}}}\otimes\mathbf{W}$. By [51, Theorem 4.2.12], the eigenvalues of ${\tilde{\mathbf{A}}}\otimes\mathbf{W}$ are $\{\lambda_{i}({\tilde{\mathbf{A}}})\,\lambda_{j}(\mathbf{W})\}_{i=1,\dots,N,\,j=1,\dots,D}$ where $\lambda_{i}({\tilde{\mathbf{A}}})$ (resp. $\lambda_{j}(\mathbf{W})$) is the $i^{th}$ (resp. $j^{th}$) eigenvalue of ${\tilde{\mathbf{A}}}$ (resp. $\mathbf{W}$). As usual, we order its eigenvalues order in a non-increasing manner [^3] Note that the order is not necessarily strictly increasing as eigenvalues may have non-trivial multiplicities.. The Cauchy-Schwarz inequality and the definition of the norm $\|\cdot\|_{2}$ imply that
$\|{\tilde{\mathbf{A}}}\otimes\mathbf{W}\|_{2}\leq\|{\tilde{\mathbf{A}}}\|_{2}\|\mathbf{W}\|_{2}=\Big{(}\sum_{i=1}^{N}\,\lambda_{i}({\tilde{\mathbf{A}}})^{2}\Big{)}^{1/2}\,\Big{(}\sum_{j=1}^{D}\,\lambda_{j}(\mathbf{W})^{2}\Big{)}^{1/2}.$ (26)
Together, equation 25 and equation 26 provide the following sufficient condition for invertibility of the $I_{ND}+{\tilde{\mathbf{A}}}\otimes\mathbf{W}$; and thus for the injectivity of $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}$
$\Big{(}\sum_{i=1}^{N}\,\lambda_{i}({\tilde{\mathbf{A}}})^{2}\Big{)}^{1/2}\,\Big{(}\sum_{j=1}^{D}\,\lambda_{j}(\mathbf{W})^{2}\Big{)}^{1/2}<1.$ (27)
Step 2 - Computing the eigenvalues of $\tilde{\mathbf{A}}$
Since we have assumed that $\mathcal{G}$ is a complete graph on $N$ nodes then $\tilde{\mathbf{A}}_{i,j}=1/N$ for each $i,j=1,\dots,N$. In particular, $\mathcal{G}$ has $N$ identical rows. Thus, 00 is an eigenvalue of $\mathcal{G}$ with multiplicity at-least $N-1$. Additionally, one easily verifies that $\mathbf{1}_{N}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(1,\dots,1)\in\mathbb{R}^{N}$ (the vector with all components equal to $1$) is an eigenvector of $\tilde{\mathbf{A}}$. It corresponds to the eigenvalue $\lambda_{1}(\tilde{\mathbf{A}})=1$ since

### v2

@@ 5 Experimental Validation
Model Removed ogbn-proteins pokec ogbn-arxiv ogbn-products SGFormer (Wu et al., 2023b) N/A 79.53 ± 0.38 73.76 ± 0.24 72.63 ± 0.13 89.09 ± 0.10 SMPNN N/A 83.15 ± 0.24 79.76 ± 0.19 73.75 ± 0.24 90.61 ± 0.05 SMPNN Residual 68.49 ± 2.59 68.17 ± 8.22 39.67 ± 14.60 89.69 ± 0.05 SMPNN $\alpha$ 82.90 ± 0.34 80.10 ± 0.12 73.00 ± 0.59 90.77 ± 0.04 SMPNN FF 80.51 ± 0.61 78.40 ± 0.20 73.25 ± 0.58 90.01 ± 0.10 SMPNN GCN LayerNorm 80.74 ± 0.91 79.42 ± 0.15 74.46 ± 0.22 90.54 ± 0.08
@@
Deep Models are Possible with SMPNNs. It is well known that conventional message-passing GNNs (Kipf and Welling, 2017; Veličković et al., 2018; Brody et al., 2021) are restricted to shallow architectures, since their performance degrades when stacking many layers. Next, we demonstrate that deep models are possible with SMPNNs. In particular, we perform an experiment in which we progressively increase the number of SMPNN layers while keeping all other hyperparameters constant for the ogbn-arxiv and ogbn-proteins datasets, from 2 to 12 layers (Table 7). Adding up to 6 layers seems to improve performance, and it plateaus thereafter. Additionally, we perform another experiment in which we follow the same procedure but for an SMPNN without residual connections after convolutions. In this case, we observe a clear drop in performance after 4 layers, which aligns with our theoretical understanding from Section 4.
@@
GPU Memory Scaling. We proceed to examine the model’s scalability utilizing the ogbn-products dataset and randomly sampling a subset of nodes for graph mini-batching (from 10K to 100K), following the analysis in (Wu et al., 2023b). In Figure 2 (Appendix F), we report the maximum GPU memory usage in GB against the number of edges in the subgraph induced by the sampled nodes. We can see that it asymptotes towards linearity in the number of edges as discussed in Section 3.2. For small subgraphs $N$ dominates the complexity term and results in a steeper slope. Additionally, we compare the computational overhead of different SMPNN configurations to that of SGFormer on a Tesla V100 GPU, see Figure 3 (Appendix F). Interestingly, the scaling factors $\alpha$ seem to have a substantial computational impact as the number of nodes increases, which, combined with the ablations in Table 6, suggests it may be better to remove them for very large graphs. Moreover, we find that the GPU consumption of SMPNNs can be substantially reduced when omitting the pointwise feedforward part of the block. This reduces computation cost substantially below that of SGFormer. Hence, one can trade GPU usage for performance: note that as shown in Table 6, SMPNNs without feedforwards still outperform the SGFormer baseline. SGFormer does not use feedforward networks or traditional attention mechanisms. Instead, it employs a simplified linear version of attention and even omits the softmax activation function.
@@ 6 Conclusion
Our SMPNN framework demonstrates that standard graph convolution, when packaged within a Pre-LN Transformer-style residual block, provides a simple yet effective architecture. We empirically show that SMPNNs enable deep message-passing networks without the degradation traditionally observed in GNNs and scale efficiently to large graphs, achieving strong empirical performance and outperforming recent scalable Graph Transformers without relying on global attention. We complement these results with a new universality-based theoretical analysis explaining why residualized convolution preserves expressivity. We do not claim to introduce a fundamentally new primitive.
Our results suggest that attention is often unnecessary (or provides only marginal gains) on “traditional” large transductive graphs. This may be partly explained by the local nature and high MaxSCC ratios of commonly used benchmarks. However, more broadly and importantly, the limited impact of attention in these settings may also arise from the absence of positional encodings, which are rarely used in large-graph models due to their computational cost (e.g., Laplacian positional encodings). Without positional signals to distinguish node positions, attention mechanisms may effectively degenerate into computing near-average feature aggregations. This consideration is likely important for emerging long-range graph benchmarks (Liang et al., 2026; Bamberger et al., 2025), which we leave for future work.
Appendix A Additional Background on Graph Representation Learning
Graph Representation Learning and Notation. We denote a graph with $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents a set of nodes (or vertices), and $\mathcal{E}\subseteq(\mathcal{V}\times\mathcal{V})$ is a set of 2-tuples signifying edges (or links) within the graph. For any pair of nodes $v_{i}$ and $v_{j}$ within $\mathcal{G}$, their connection is encoded as $(v_{i},v_{j})\in\mathcal{E}$ if the edge originates from $v_{i}$ and terminates at $v_{j}$. The (one-hop) neighborhood of node $v_{i}$ constitutes the set of nodes sharing an edge with $v_{i}$, denoted by $\mathcal{N}(v_{i})=\{v_{j}|(v_{i},v_{j})\in\mathcal{E}\}$. To encode the structural connectivity amongst nodes in a graph of $N=|\mathcal{V}|$ nodes and $E=|\mathcal{E}|$ edges, an adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ is employed. This adjacency matrix may adopt a weighted or unweighted representation. In the case of a weighted adjacency matrix, the entries $A_{ij}\in\mathbb{R}$ symbolize the strength of the connection, with $A_{ij}=0$ denoting absence of a connection if $(v_{i},v_{j})\notin\mathcal{E}$. Conversely, in an unweighted adjacency matrix, $A_{ij}=1$ signifies the presence of an edge, and $A_{ij}=0$ otherwise. The degree matrix $\mathbf{D}\in\mathbb{R}^{N\times N}$ is defined as the matrix where each entry on the diagonal is the row-sum of the adjacency matrix: $D_{ii}=\sum_{j}A_{ij}$. Based on this, we can define the graph Laplacian as $\mathbf{L}=\mathbf{D}-\mathbf{A}$. The normalized graph Laplacian, denoted as $\mathbf{L}_{\text{norm}}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$, is a variation of the graph Laplacian that takes into account the degrees of the nodes. Additionally, within our context, we are interested in graphs with node features. Each node $v_{i}\in\mathcal{V}$ is accompanied by a $D$-dimensional feature vector $\mathbf{x}_{i}\in\mathbb{R}^{1\times D}$. The feature vectors for all nodes within the graph can be aggregated into a single matrix $\mathbf{X}\in\mathbb{R}^{N\times D}$ by stacking them along the row dimension.
Message Passing Neural Networks (MPNNs) are a class of Graph Neural Networks (GNNs). MPNNs operate by means of message passing, wherein nodes exchange vector-based messages to refine their representations, leveraging the graph connectivity structure as a geometric prior. Following Bronstein et al. (2021), a message passing GNN layer $l$ (excluding edge and graph-level features for simplicity) over a graph $\mathcal{G}$ is defined as: $\mathbf{x}_{i}^{(l+1)}=\phi\biggl(\mathbf{x}_{i}^{(l)},\bigoplus_{j\in\mathcal{N}(v_{i})}\psi(\mathbf{x}_{i}^{(l)},\mathbf{x}_{j}^{(l)})\biggr),$ where $\psi$ denotes a message passing function, $\bigoplus$ is some permutation-invariant aggregation operator, and $\phi$ is a readout or update function. Both $\psi$ and $\phi$ are learnable and typically implemented as MLPs. It is important to note that the update equation is local, and that at each layer each node only communicates with its one-hop neighborhood given by the adjacency matrix of the graph. More distant nodes are accessed by stacking multiple message-passing layers.
Appendix B Augmenting SMPNNs with Attention
@@
Linear Global Attention. Global attention can be seen as computing message passing over a fully connected graph. In our block, attention is computed with respect to a virtual node rather than through full $\mathcal{O}(N^{2})$ pairwise attention, which is more suitable for scaling to large graphs. Given the input query, key, and value matrices,
$\mathbf{Q}=\mathbf{X}^{(l)}\mathbf{W}_{Q}\in\mathbb{R}^{N\times D},\qquad\mathbf{K}=\mathbf{X}^{(l)}\mathbf{W}_{K}\in\mathbb{R}^{N\times D},\qquad\mathbf{V}=\mathbf{X}^{(l)}\mathbf{W}_{V}\in\mathbb{R}^{N\times D},$
we apply the following operations. First, we compute a summed query vector incorporating contributions from all nodes, and normalize both this summed query vector and the keys:
$\mathbf{Q}_{\mathrm{sn}}=\frac{\sum_{i=1}^{N}\mathbf{Q}_{i}}{\left\|\sum_{i=1}^{N}\mathbf{Q}_{i}\right\|_{2}}\in\mathbb{R}^{1\times D},\qquad\mathbf{K}_{n,i}=\frac{\mathbf{K}_{i}}{\|\mathbf{K}_{i}\|_{2}}\in\mathbb{R}^{1\times D},$ (8)
for each node $i=1,\dots,N$. Based on these, we compute attention weights
$\mathbf{A}_{\mathrm{attention}}=\mathrm{softmax}\!\left(\mathbf{Q}_{\mathrm{sn}}\mathbf{K}_{n}^{T}\right)\in\mathbb{R}^{1\times N},$ (9)
where $\mathbf{K}_{n}\in\mathbb{R}^{N\times D}$ is the matrix whose $i$-th row is $\mathbf{K}_{n,i}$. We then use these weights to aggregate the values into a single global feature vector
$\mathbf{g}^{(l)}=\mathbf{A}_{\mathrm{attention}}\mathbf{V}\in\mathbb{R}^{1\times D},$ (10)
which is broadcast to all nodes:
$\mathbf{X}^{(l+1)}=\mathbf{1}_{N}\mathbf{g}^{(l)}\in\mathbb{R}^{N\times D},$ (11)
where $\mathbf{1}_{N}\in\mathbb{R}^{N\times 1}$ is a column vector of ones. Equivalently,
$\mathbf{X}^{(l+1)}=\left(\mathbf{1}_{N}\mathbf{A}_{\mathrm{attention}}\right)\mathbf{V},$ (12)
where $\mathbf{1}_{N}\mathbf{A}_{\mathrm{attention}}\in\mathbb{R}^{N\times N}$ is a matrix with all rows equal to the same attention vector.
The operations above extend naturally to multi-head attention. If we use $H$ attention heads, then for each head $h$ we compute
$\mathbf{A}_{\mathrm{attention}}^{(h)}=\mathrm{softmax}\!\left(\mathbf{Q}_{\mathrm{sn}}^{(h)}(\mathbf{K}_{n}^{(h)})^{T}\right)\in\mathbb{R}^{1\times N},$ (13)
and the corresponding head output
$\mathbf{g}^{(h)}=\mathbf{A}_{\mathrm{attention}}^{(h)}\mathbf{V}^{(h)}\in\mathbb{R}^{1\times D}.$ (14)
The final global feature matrix is then obtained by aggregating (using the mean) the contributions from all heads:
$\mathbf{X}^{(l+1)}=\mathbf{1}_{N}\left(\frac{1}{H}\sum_{h=1}^{H}\mathbf{g}^{(h)}\right)\in\mathbb{R}^{N\times D}.$ (15)
@@
The update equations above are reminiscent of hybrid Graph Transformers (Rampášek et al., 2022; Wu et al., 2022; 2023a; 2023b), and we find that they do not provide tangible improvements over standard SMPNNs. Note that we also experimented with other attention mechanisms, which did not help.
Appendix C Oversmoothing and Residual Connections
Oversmoothing is regarded as the tendency of node features to approach the same value after several message-passing transformations. This phenomenon has prompted the adoption of relatively shallow GNNs, as adding many layers has traditionally resulted in node-level features that are too similar to each other and, hence, indistinguishable for downstream learners. This has hindered scalability and led to models with orders of magnitude fewer parameters than, for instance, counterparts in the literature on 2D generative modeling (Esser et al., 2024) and LLMs (Touvron et al., 2023), which have already adopted architectures with billions of parameters.
In general, standalone message-passing GNNs tend to behave as low-pass filters and may consequently lead to oversmoothing. However, this does not always need to be the case: graph convolution can also magnify high frequencies if it is augmented with residual connections. Following (Giovanni et al., 2023), message-passing is characterized as being low-frequency dominant (LFD) or high-frequency dominant (HFD) depending on the asymptotic behavior of the normalized Dirichlet energy of the system as the number of diffusion layers $l\rightarrow\infty$. These lead to oversmoothing and oversharpening, respectively.
Definition C.1 (Graph Dirichlet Energy of a Message Passing System (Zhou and Scholkopf, 2005)).
@@ Appendix C Oversmoothing and Residual Connections
Definition C.2 (Graph Frequency Dominance (Giovanni et al., 2023)).